OpenStudy (anonymous):

Use the simple interest formula to find the missing value. Rates are annual rates. Assume 360 days per year and 30 days per month. p = ?, r = 4.5%, t = 10 days, i = $26.68 A. p = $21,344.00 B. p = $711.47 C. p = $7114.67 D. p = $21,670.09

OpenStudy (experimentx):

use r=4.5/360 I = P((1+r/100)^10 - 1)

OpenStudy (experimentx):

use the above relation

OpenStudy (anonymous):

A. p = $21,344.00